OpenStudy (anonymous):

find the solution w(9+w)=0 a) 0 & -9 b) 0 & 9 c) 0 & 81

OpenStudy (anonymous):

w(w+9) =0 will be satisfied only when w = 0 or w + 9 = 0 w = 0, -9